Suppose the Acme Drug Company develops a new drug, designed to prevent colds. The company


states that the drug is equally effective for men and women. To test this claim, they choose a


simple random sample of 110 women and 190 men from a population of 100,000 volunteers.


At the end of the study, 28% of the women caught a cold; and 61% of the men caught a cold.


Based on these findings, can we reject the company's claim that the drug is equally effective for


men and women? Use a 0.05 level of significance.

1
Expert's answer
2022-06-13T23:25:31-0400

The value of the pooled proportion is computed as


"\bar{x}=\dfrac{X_1+X_2}{n_1+n_2}=\dfrac{\hat{p}_1n_1+\hat{p}_2n_2}{n_1+n_2}"

"=\dfrac{0.28(110)+0.61(190)}{110+190}=0.489"

The following null and alternative hypotheses need to be tested:

"H_0:p_1=p_2"

"H_1:p_1\not=p_1"

This corresponds to a two-tailed test, and a z-test for two population proportions will be used.

Based on the information provided, the significance level is "\alpha = 0.05," and the critical value for a two-tailed test is "z_c = 1.96."

The rejection region for this two-tailed test is "R = \{z:|z|>1.96\}."

The z-statistic is computed as follows:


"z=\dfrac{\hat{p}_1-\hat{p}_2}{\sqrt{\bar{p}(1-\bar{p})(1/n_1+1/n_2)}}"


"=\dfrac{0.28-0.61}{\sqrt{0.489(1-0.489)(1/110+1/190)}}=-5.5101"

Since it is observed that "|z|=5.5101>1.96=z_c," it is then concluded that the null hypothesis is rejected.

Using the P-value approach:

The p-value is "p=2P(z<-5.5101)=0," and since "p=0<0.05=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population proportion "p_1" is different than "p_2," at the "\alpha = 0.05" significance level.